OpenStudy (anonymous):

Which functions have a positive rate of change and which have a negative rate of change?

OpenStudy (anonymous):

A)positive B)negative 1)2x-5y=-15 ------------- 2.y = −4x + 9 ------------- 3)x y 2 –10 4 –16 6 –22 __________ 4) x y –3 15 –2 11 –1 7

OpenStudy (anonymous):

no not mind thr fffd things

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@Snowmancer

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

looks like they are all going down, except for the first one which will go up

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@misty1212 plz come back

OpenStudy (misty1212):

for 2, 3, 4 as x gets bigger, y gets smaller

OpenStudy (misty1212):

for \(2x-5y=-15\) the slopes is\(\frac{2}{5}\)so as x gets bigger, so does y

OpenStudy (anonymous):

ok 2,3,4 are negitive and 1 positive

OpenStudy (anonymous):

plzz i am begging @misty1212